# spacing=nonfrench

### v1

@@ 3.3 Boundary Values of Functions in $\mathcal{B}_{p,p}^{s}(\Omega)$
Let $\Gamma$ be a quasicircle and $\Omega$ one of its complementary domains $\Omega_{i,e}$ as above. The boundary values of the critical case $s=1/p$ have been treated in the last section. As announced in the introduction, we will define now what we mean by boundary values of functions in $\mathcal{B}_{p,p}^{s}(\Omega),\,p>1,\,s\in(0,1)$. As we shall see the cases $s<1/p$ and $s>1/p$ will show to be very different. We start with the case of $s>1/p$.
@@
Let $\Omega$ be a (bounded or unbounded) domain bounded by a bounded Jordan curve $\Gamma$. For any $u\in\mathcal{B}_{p,p}^{s}(\Omega)$ with $s\in(1/p,1)$ there exists a constant $C>0$ such that
@@
Let $z\in\Omega$ and $D$ is the disk $D(z,d(z,\Gamma)/2)$. By the mean value property of harmonic functions one may write
@@
(Here and in the sequel we write $|D|$ for the Lebesgue measure of $D$.) so that, by Hölder inequality,
@@
Let $\Omega$ be a (bounded or unbounded) domain bounded by a quasicircle $\Gamma$ and $\alpha\in(0,1]$. The following are equivalent:
@@
If now $s\in(1/p,1)$ then $\alpha=s-1/p\in(0,1)$ and Gehring-Martio theorem implies that functions in $\mathcal{B}_{p,p}^{s}(\Omega)$ are continuous on $\overline{\Omega}$ , so that boundary values of functions in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ are well-defined and they characterize $u$, as being the unique harmonic extension of this continuous boundary value. We may now define the space $\mathcal{B}_{p,p}^{s}(\Gamma)$ for this range $s\in(1/p,1)$ as being the space of (s1/p)(s-1/p)-Hölder functions $f$ on $\Gamma$ whose harmonic extension $u_{i,e}$ to $\Omega_{i,e}$ belongs to $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$, and the space $\mathcal{B}_{p,p}^{s}(\Gamma)$ is assigned the natural norm $\|\cdot\|_{\mathcal{B}_{p,p}^{s}(\Gamma)}$ so that $\|f\|^{p}_{\mathcal{B}_{p,p}^{s}(\Gamma)}=\|u_{i}\|^{p}_{\mathcal{B}_{p,p}^{s}(\Omega_{i})}+\|u_{e}\|^{p}_{\mathcal{B}_{p,p}^{s}(\Omega_{e})}$.
3.3.1 Trace Mappings
As we have already seen, in the case of $s>1/p$ the spaces $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ have well-defined boundary values and, with the terminology adopted in the introduction, the trace mapping
@@ 3.3.1 Trace Mappings
is an isomorphism. The first purpose of this part is to determine conditions that imply
@@
We already know the result for $s=1/p$ and $h(\Gamma)<p<h(\Gamma)/(1-h(\Gamma))$ (but in this case the boundary value has a different meaning). Now the space $\Lambda^{s}(\Omega)$ introduced above in connection with Gehring-Martio theorem coincides with the Besov space $\mathcal{B}_{\infty,\infty}^{s}(\Omega)$. A theorem of Hinkkanen [20] (see also [48]) states that suppose the curve $\Gamma$ is a Jordan curve and if $f\in\Lambda^{s}(\Gamma)$ with $s<1/2$ then the harmonic extension of $f$ to $\Omega_{i,e}$ is in $\mathcal{B}_{\infty,\infty}^{s}(\Omega)$ with norm bounded by that of $f$ and the index $1/2$ is critical. Hinkkanen’s result is true for all Jordan curves, and it has been improved to $s<1/K$ in [48] for quasicircles such that their Riemann mapping has a $K$-quasiconformal extension with $1\leq K\leq 2$. We thus define
@@
We may now state the main result of this part.
Theorem 3.12.
Let $\Gamma,p,s$ be such that
$h(\Gamma)<p\leq\infty,\;1/p\leq s\;\mathrm{and}\;s<\frac{p+K_{*}-h(\Gamma)}{K_{*}p}$ (3.6)
then
$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma).$ (3.7)
Proof.
Let $P_{e}$, as before, be the operator which takes a function on $\Gamma$ to its harmonic extension in $\Omega_{e}$. As we have recalled that if
@@
are both bounded. We want to interpolate these results. Using complex interpolation, we have, for $0<\theta<1$,
$\mathcal{B}_{q,q}^{1/q}(\Omega),\mathcal{B}_{\infty,\infty}^{\sigma}(\Omega)]_{\theta}=\mathcal{B}_{p,p}^{s}(\Omega),$
@@
We must decide what the couples $(p,s)$ are that may be "caught" in this way in view of the constraints: the relations above are equivalent to
$\sigma$ $=\frac{s-\frac{1}{p}}{\theta},$
$q$ $=(1-\theta)p,$
and the constraints are
$0<\frac{s-\frac{1}{p}}{\theta}$ $<1/K_{*},$
$h(\Gamma)<(1-\theta)p$ $<\frac{h(\Gamma)}{h(\Gamma)-1}.$
That correspond to $\theta$ being in the intersection of two intervals
$(K_{*}(s-\frac{1}{p}),1),\;(1-\frac{h(\Gamma)}{p(h(\Gamma)-1)},1-\frac{h(\Gamma)}{p}).$
We see that in order to have a solution we must have $s>1/p,p>h(\Gamma)$, and $K_{*}(s-\frac{1}{p})<1-\frac{h(\Gamma)}{p}$, which is the condition of the statement of the theorem. ∎
3.3.2 Weighted Sobolev Spaces
Let $\omega$ be an $A_{p}$ weight defined in $\mathbb{C}$, for some $p>1$. The weighted Sobolev space $W^{1,p}(\omega,\mathbb{C})$ is the space of tempered distributions $T$ such that $\nabla T$, taken in the sense of distributions, is a function in $L^{p}(\omega,\mathbb{C})$. We will use these spaces for $\omega(z)=d(z,\Gamma)^{(1-s)p-1}\in A_{p}$ provided that $(h(\Gamma),p,s)\in\mathcal{A}$.
For the usual technical reasons we need to know that $C_{c}^{\infty}(\mathbb{C})$-functions are dense in $W^{1,p}(\omega,\mathbb{C})$.
@@ 3.3.2 Weighted Sobolev Spaces
Let us consider $f\in W^{1,p}(\omega,\mathbb{C})$: without loss of generally we may assume that $f$ has compact support. Let then $(\varphi_{\varepsilon})$ be a $C_{c}^{\infty}(\mathbb{C})$-approximation of identity, so that $f_{\varepsilon}=f*\varphi_{\varepsilon}\in C_{c}^{\infty}(\mathbb{C})$. Now
@@
which converges to 0 by the Lebesgue dominated convergence theorem. Indeed $|\nabla f_{\varepsilon}|$ is controlled by the Hardy-Littlewood maximal function of $\nabla f$ and this maximal function is in $L^{p}(\omega,\mathbb{C})$ from the fact that $\omega\in A_{p}$. ∎
In the classical case $\omega\equiv 1$ we recall that if $p>2$ then $W^{1,p}(\mathbb{C})\subset\Lambda^{\alpha}(\mathbb{C})$ with $\alpha=1-2/p$. In the weighted case we do not know this embedding result in general but we have two particular analogues.
@@
The following two embedding assertions hold:
  • If $\omega\in A_{1}$ then every function in $W^{1,p}(\omega,\mathbb{C})$, $p>2$, is continuous;
  • If $\omega=d(z,\Gamma)^{(1-s)p-1}$ is only assumed to be in $A_{q}$ for some $q>1$ then the same result is true if $p>\frac{6-2h(\Gamma)}{2-h(\Gamma)}$.
Remark 3.15.
When $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$ it is known that for any $q>1$, $\omega\in A_{q}\Leftrightarrow\omega\in A_{1}$ if 0<(1s)p<10<(1-s)p<1, i.e., s>11/ps>1-1/p.
@@
Suppose first that $\omega\in A_{1}$. Choose $q$ so that $2<q<p$ and let $B$ be an open ball in $\mathbb{C}$ . By Hölder inequality, we get
$\int_{B}|\nabla f(x)|^{q}dxdy\leq\left(\int_{B}|\nabla f(x)|^{p}\omega(x)dxdy\right)^{\frac{q}{p}}\left(\int_{B}\omega(x)^{-\frac{q}{p-q}}dxdy\right)^{1-\frac{q}{p}}<\infty$
because $f\in W^{1,p}(\omega,\mathbb{C})$, and $\omega\in A_{1}\subset A_{p/q}$, which is equivalent to $\omega^{-q/(p-q)}\in A_{p/(p-q)}$, and in particular implies that $\omega^{-q/(p-q)}$ is locally integrable.
For the second case let us recall first that if $B$ is a ball containing $\Gamma$ then
$\iint_{B}d(z,\Gamma)^{\beta}dxdy<\infty$
if $\beta>h(\Gamma)-2$. By the previous computation it follows that a sufficient condition for the validity of the theorem is
$\frac{2}{p-2}>h(\Gamma)-2$
which is equivalent to the condition stated. ∎
We return to Theorem 3.12. It states that under certain hypothesis on $(h(\Gamma),p,s)$ we have
Let us call this space of boundary values $\mathcal{B}_{p,p}^{s}(\Gamma)$.
@@
With the following assumption
$\frac{1}{p}<1-\frac{1}{p}\leq s<\frac{p+K_{*}-h(\Gamma)}{K_{*}p}$ (3.8)
we have that the elements of $W^{1,p}(\omega,\mathbb{C})$ are continuous functions in $\mathbb{C}$ and that
@@
where $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$ and $W^{1,p}(\omega,\mathbb{C})|_{\Gamma}=\{f|_{\Gamma}:\,f\in W^{1,p}(\omega,\mathbb{C})\}.$
Before we start the proof, let us comment on this statement. First of all, the assumption made implies that $\omega\in A_{p}$ and that
$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma)\;(=\mathcal{B}_{p,p}^{s}(\Gamma)).$
Moreover the inequality 1/p<11/p1/p<1-1/p is equivalent to $p>2$ and the inequality s>11/ps>1-1/p further implies, as we have noticed above, that $\omega\in A_{1}$, from which it follows that the elements of $W^{1,p}(\omega,\mathbb{C})$ are continuous functions (Proposition 3.14). Finally, the inequality
$1-\frac{1}{p}<\frac{p+K_{*}-h(\Gamma)}{K_{*}p}$
is equivalent to $p<\frac{2K_{*}-h(\Gamma)}{K_{*}-1}$.
@@
We first show that $\mathcal{B}_{p,p}^{s}(\Gamma)\subset W^{1,p}(\omega,\mathbb{C})|\Gamma$. Let $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$ and let $u_{i}$ the harmonic extension of $f$ to $\Omega_{i}$ . By hypothesis $u_{i}\in W^{1,p}(\omega,\Omega_{i})$ and a theorem of [8] asserts that since $\Gamma$ is a quasicircle and $\omega\in A_{p}$, $u_{i}$ can be extended to a function of $W^{1,p}(\omega,\mathbb{C})$.
@@
are holomorphic respectively in $\Omega_{i}$ and $\Omega_{e}$, and the fact that $\omega\in A_{p}$ implies that $F_{i,e}\in\mathcal{B}_{p,p}^{s}(\Omega_{i,e}\to\Gamma)$, which is equal to $\mathcal{B}_{p,p}^{s}(\Gamma)$ under the assumption of Theorem 3.12. Using $f|_{\Gamma}=F_{i}-F_{e}$ we may see $f|_{\Gamma}\in\mathcal{B}_{p,p}^{s}(\Gamma)$.
@@
Under the assumption of Theorem 3.16, the space $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$ is dense in $\mathcal{B}_{p,p}^{s}(\Gamma)$, and thus the conclusions in Theorem 3.5 hold for all $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$.
We end this section by a question. Is it true that within the hypothesis of Theorem 3.16 if $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$ then its extension to the whole plane by the harmonic extension on both sides belongs to $W^{1,p}(\omega,\mathbb{C})$? We do not know the answer to this question in general but only for the cases with improved regularity (bigger $s$):
@@
Let $\Gamma$ be a quasicircle and let $f$ be a function which is $\alpha$-Hölder continuous in $\mathbb{C}$ with $\alpha>h(\Gamma)-1$ and of class $C^{1}$ in $\mathbb{C}\backslash\Gamma$. Then if the function $F(z)=\nabla f(z),\,z\in\mathbb{C}\backslash\Gamma$ (thus defined almost everywhere) is in $L^{1}_{loc}(\mathbb{C})$, we have that, in the sense of distributions, $\nabla f=F$.
@@
Let $\varphi$ be a $C_{c}^{\infty}$-function in $\mathbb{R}$ with support included in [1,1][-1,1]. As usual we define the approximation of identity $\varphi_{\varepsilon}(x)=\frac{1}{\varepsilon}\varphi(\frac{x}{\varepsilon})$. From this we can define the approximation of identity in two variables
@@
We may assume without loss of generality that $f$ has compact support, the function $f*\psi_{\varepsilon}$ is then a $C^{\infty}$-function with compact support and it converges uniformly in $\mathbb{C}$ towards $f$. From this we deduce that
@@
We split the left-hand side into two parts:
Let $\Gamma_{\varepsilon}$ be the set of points at distance less than $2\varepsilon$ from $\Gamma$. Then
@@
because $\int_{[-\varepsilon,\varepsilon]}\varphi_{\varepsilon}^{\prime}(t_{1})dt_{1}=0$. Now $|f(x-t_{1},y-t_{2})-f(x,y-t_{2})|\leq C\varepsilon^{\alpha}$ and $|\varphi_{\varepsilon}^{\prime}(t_{1})|\leq C\varepsilon^{-2}$. Since we integrate on an interval of length $2\varepsilon$, the absolute value of the whole integral is bounded by $C\varepsilon^{\alpha-1}$.
@@
But we know that $|K\Gamma_{\varepsilon}|\varepsilon^{h(\Gamma)-2}\leq C$; we conclude that
@@
Based on Proposition 3.18, a sufficient condition for a positive answer to the question is the hypothesis of Theorem 3.16 and $s>\frac{1}{p}+(h(\Gamma)-1)$. To be precise, the function $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$, defined to be $u_{i,e}$ in $\Omega_{i,e}$, satisfies the assumption of Proposition 3.18, and thus belongs to $L^{1}_{\text{loc}}(\mathbb{C})$ in the sense of distribution, so that $f\in W^{1,p}(\omega,\mathbb{C})$. But for the condition $s>\frac{1}{p}+(h(\Gamma)-1)$ to be compatible with (3.8) some elementary arithmetic left to the reader leads to the following sufficient conditions:
$p>2,\quad\frac{1}{p}+h(\Gamma)-1<s<\frac{p+K_{*}-h(\Gamma)}{pK_{*}}.$
Notice that in the case of a chord-arc curve ($h(\Gamma)=1$) there are admissible values of $s$ for all $p>2$. In general, for having amissible values of $s$ one needs
$h(\Gamma)<\frac{K_{*}+1}{K_{*}+\frac{1}{p}}$
and for this inequality to hold for any $p>2$ and $K_{*}\leq 2$, $h(\Gamma)<\frac{6}{5}$ is necessary and sufficient.
3.3.3 Almost-Dirichlet Principle
We start by recalling the classical Dirichlet principle. If $\Omega$ is a Jordan domain with $\partial\Omega=\Gamma$ and $f$ is a continuous function on $\Gamma$ having a continuous extension $F:\overline{\Omega}\to\mathbb{C}$ which is in $C^{1}(\Omega)$. then, if $u$ stands for the harmonic extension of $f$ to $\Omega$, we have

### v2

@@ 3.1 Boundary values of functions in $\mathcal{B}_{p,p}^{s}(\Omega)$
Let $\Gamma$ be a quasicircle and $\Omega$ one of its complementary domains $\Omega_{i,e}$ as above. The boundary values of $\mathcal{B}_{p,p}^{s}(\Omega)$-functions in the critical case $s=1/p$ have been treated in the last section. As announced in the introduction, we will define now what we mean by the boundary values of the functions in $\mathcal{B}_{p,p}^{s}(\Omega),\,p>1,\,s\in(0,1)$. As we shall see, the cases $s<1/p$ and $s>1/p$ will show to be very different. We now consider the case of $s>1/p$.
@@
Let $\Omega$ be a domain bounded by a bounded Jordan curve $\Gamma$. For any $u\in\mathcal{B}_{p,p}^{s}(\Omega)$ with $s\in(1/p,1)$ there exists a constant $C>0$ such that
@@
Let $z\in\Omega$ and $D$ be the disk $D(z,d(z,\Gamma)/2)$. By the mean value property of harmonic functions, one may write
@@
so that, by Hölder inequality,
@@
Let $\Omega$ be a domain bounded by a quasicircle $\Gamma$ and $\alpha\in(0,1]$. The following are equivalent:
@@
If now $s\in(1/p,1)$ then $\alpha=s-1/p\in(0,1)$ and Gehring-Martio theorem implies that the functions in $\mathcal{B}_{p,p}^{s}(\Omega)$ are continuous on $\overline{\Omega}$ , so the boundary values of the functions in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ are well-defined and characterize $u$, as being the unique harmonic extension of this continuous boundary value. We may now define the space $\mathcal{B}_{p,p}^{s}(\Gamma)$ for this range $s\in(1/p,1)$ as being the space of (s1/p)(s-1/p)-Hölder functions $f$ on $\Gamma$ whose harmonic extension $u_{i,e}$ to $\Omega_{i,e}$ belongs to $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$, and the space $\mathcal{B}_{p,p}^{s}(\Gamma)$ is assigned the natural norm $\|\cdot\|_{\mathcal{B}_{p,p}^{s}(\Gamma)}$ so that $\|f\|^{p}_{\mathcal{B}_{p,p}^{s}(\Gamma)}=\|u_{i}\|^{p}_{\mathcal{B}_{p,p}^{s}(\Omega_{i})}+\|u_{e}\|^{p}_{\mathcal{B}_{p,p}^{s}(\Omega_{e})}$.
3.2 Trace mappings on $\mathcal{B}_{p,p}^{s}(\Omega)$
As we have already seen, in the case $s>1/p$ the spaces $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ have well-defined boundary values and, with the terminology adopted in the introduction, the trace mapping
@@ 3.2 Trace mappings on $\mathcal{B}_{p,p}^{s}(\Omega)$
is an isomorphism. The goal of this subsection is to discuss conditions that imply
@@
We already know the result for $s=1/p$ and $h(\Gamma)<p<h(\Gamma)/(1-h(\Gamma))$ (but in this case the boundary value has a different meaning).
Another case is $p=\infty$: the space $\Lambda^{s}(\Omega)$ introduced above in connection with the Gehring-Martio theorem coincides with the Besov space $\mathcal{B}_{\infty,\infty}^{s}(\Omega)$. A theorem of Hinkkanen [20] (see also [54]) states that if the curve $\Gamma$ is any Jordan curve and if $f\in\Lambda^{s}(\Gamma)$ with $s<1/2$ then the harmonic extension of $f$ to $\Omega_{i,e}$ is in $\mathcal{B}_{\infty,\infty}^{s}(\Omega)$ with norm bounded by that of $f$ and the index $1/2$ is critical. Hinkkanen’s result is true for all Jordan curves, and it has been improved to $s<1/K$ in [54] for quasicircles such that their Riemann mapping has a $K$-quasiconformal extension with $1\leq K\leq 2$. Thus, we define
@@
Let $P_{e}$, as before, be the operator that takes a function on $\Gamma$ to its harmonic extension in $\Omega_{e}$. We have recalled that if
@@
are both bounded. We want to interpolate these results. It is known that in the sense of complex interpolation we have in the case of unit disk $\mathbb{D}$, for $0<\theta<1$,
$\mathcal{B}_{q,q}^{1/q}(\mathbb{D}),\mathcal{B}_{\infty,\infty}^{\sigma}(\mathbb{D})]_{\theta}=\mathcal{B}_{p,p}^{s}(\mathbb{D}),$
@@
If the interpolation statement above remained true for quasicircles then, using interpolation for the operator $P_{e}$, we would produce many new examples of equality of Besov-Sobolev spaces from inside and outside $\Gamma$. We strongly believe that this interpolation for quasicircles is true but we could not prove it so far, except for the rectifiable case to be treated in Section 4.
3.3 Trace mappings on weighted Sobolev spaces
Let $\omega$ be an $A_{p}$-weight defined in $\mathbb{C}$, for some $p>1$. The weighted Sobolev space $W^{1,p}(\omega,\mathbb{C})$ is the space of tempered distributions $T$ such that $T$ is locally integrable and $\nabla T$, taken in the sense of distributions, is a function in $L^{p}(\omega,\mathbb{C})$. We will use these spaces for $\omega(z)=d(z,\Gamma)^{(1-s)p-1}\in A_{p}$ provided that $(h(\Gamma),p,s)\in\mathcal{A}$.
For the usual technical reasons, we need to know that the $C_{c}^{\infty}(\mathbb{C})$-functions are dense in $W^{1,p}(\omega,\mathbb{C})$. See [8] for a proof of a more general case.
@@ 3.3 Trace mappings on weighted Sobolev spaces
Consider $f\in W^{1,p}(\omega,\mathbb{C})$: without loss of generality, we may assume that $f$ has compact support. Let then $(\varphi_{\varepsilon})$ be a $C_{c}^{\infty}(\mathbb{C})$-approximation of identity, so that $f_{\varepsilon}=f*\varphi_{\varepsilon}\in C_{c}^{\infty}(\mathbb{C})$. Now
@@
which converges to 0 by the Lebesgue dominated convergence theorem. Indeed $|\nabla f_{\varepsilon}|$ is controlled by the Hardy-Littlewood maximal function of $\nabla f$ and this maximal function is in $L^{p}(\omega,\mathbb{C})$ from the fact that $\omega\in A_{p}$ (see [34]). ∎
In the classical case $\omega\equiv 1$ we recall that if $p>2$ then $W^{1,p}(\mathbb{C})\subset\Lambda^{\alpha}(\mathbb{C})$ with $\alpha=1-2/p$, and moreover, the property being local, a similar inclusion holds for $W^{1,p}(D)$, $D$ being a disk.
@@
If $p>2$, the following two embedding assertions hold:
  • If $\omega\in A_{1}$ then all functions in $W^{1,p}(\omega,\mathbb{C})$ are continuous;
  • If $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$ is only assumed to be locally integrable, then the same result holds if $s>\frac{h(\Gamma)}{2}-\frac{h(\Gamma)-1}{p}$.
We recall here that when $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$ it is known that for any $q>1$, $\omega\in A_{q}\Leftrightarrow\omega\in A_{1}$ if $h(\Gamma)-1<(1-s)p<1$, i.e., $1-\frac{h-1}{p}s>1-1/p$.
@@
Suppose first that $\omega\in A_{1}$. Choose $q$ so that $2<q<p$ and let $D$ be an open disk in $\mathbb{C}$ . By Hölder inequality, we get
$\int_{D}|\nabla f(x)|^{q}dxdy\leq\left(\int_{D}|\nabla f(x)|^{p}\omega(x)dxdy\right)^{\frac{q}{p}}\left(\int_{D}\omega(x)^{-\frac{q}{p-q}}dxdy\right)^{1-\frac{q}{p}}<\infty$ (3.1)
because $f\in W^{1,p}(\omega,\mathbb{C})$, and $\omega\in A_{1}\subset A_{p/q}$, which is equivalent to $\omega^{-q/(p-q)}\in A_{p/(p-q)}$, and in particular implies that $\omega^{-q/(p-q)}$ is locally integrable. We conclude with the Sobolev embedding theorem mentioned above.
For the second case, if $D$ is a disk containing $\Gamma$ then by the same computation as that in Remark 2.10 we see
$\iint_{D}d(z,\Gamma)^{\beta}dxdy<\infty$
if $\beta>h(\Gamma)-2$. By the previous computation it follows that a sufficient condition for the validity of the theorem is that the second integral in the right hand-side of (3.1) is finite, and this will be the case if there exists $q\in(2,p)$ such that
$\frac{q}{p-q}((1-s)p-1)>h(\Gamma)-2.$
This condition is equivalent to, with $h=h(\Gamma)$:
F(q,s,p,h)=q(h1(1s)p)+p(2h)>0.F(q,s,p,h)=q(h-1-(1-s)p)+p(2-h)>0.
The function $F$ being affine in $q$, the last inequality occurs if and only if
$\max(F(2,s,p,h),F(p,s,p,h))>0.$
But
$F(2,s,p,h)$ $=2p\left(s-\left(\frac{h}{2}-\frac{h-1}{p}\right)\right),$
$F(p,s,p,h)$ $=p^{2}\left(s-\left(1-\frac{1}{p}\right)\right).$
The proposition easily follows by noticing that, if $p>2$,
$\frac{h}{2}-\frac{h-1}{p}<1-\frac{1}{p}.$
Let us now assume that the quasicircle $\Gamma$ satisfies
$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma),$
and we then call this space of boundary values $\mathcal{B}_{p,p}^{s}(\Gamma)$.
@@
If $p>2,\,s>\frac{h(\Gamma)}{2}-\frac{h(\Gamma)-1}{p}$ and $\omega\in A_{p}$ where $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$ then the elements of $W^{1,p}(\omega,\mathbb{C})$ are continuous functions in $\mathbb{C}$ and
@@
where $W^{1,p}(\omega,\mathbb{C})|_{\Gamma}=\{f|_{\Gamma}:\,f\in W^{1,p}(\omega,\mathbb{C})\}.$
@@
We first show that $\mathcal{B}_{p,p}^{s}(\Gamma)\subset W^{1,p}(\omega,\mathbb{C})|\Gamma$. Let $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$ and let $u_{i}$ the harmonic extension of $f$ to $\Omega_{i}$ . By hypotheses $u_{i}\in W^{1,p}(\omega,\Omega_{i})$ and a theorem of [8] asserts that since $\Gamma$ is a quasicircle and $\omega\in A_{p}$, $u_{i}$ can be extended to a function of $W^{1,p}(\omega,\mathbb{C})$.
@@
are holomorphic respectively in $\Omega_{i}$ and $\Omega_{e}$, and the fact that $\omega\in A_{p}$ implies that $F_{i,e}\in\mathcal{B}_{p,p}^{s}(\Omega_{i,e}\to\Gamma)$, which is equal to $\mathcal{B}_{p,p}^{s}(\Gamma)$ by hypotheses. Using $f|_{\Gamma}=F_{i}-F_{e}$ we may see $f|_{\Gamma}\in\mathcal{B}_{p,p}^{s}(\Gamma)$.
@@
Under the assumption of Theorem 3.5, the space $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$ is dense in $\mathcal{B}_{p,p}^{s}(\Gamma)$, and thus the conclusions of Theorem 2.9 hold for all $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$.
We end this subsection with a question. Is it true that within the hypotheses of Theorem 3.5 if $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$ then its extension to the whole plane by the harmonic extension on both sides belongs to $W^{1,p}(\omega,\mathbb{C})$? We do not know the answer to this question in general, but only for the cases with improved regularity (larger $s$):
@@
Let $\Gamma$ be a quasicircle, and let $f$ be a function that is $\alpha$-Hölder continuous in $\mathbb{C}$ with $\alpha>h(\Gamma)-1$ and of class $C^{1}$ in $\mathbb{C}\backslash\Gamma$. Then if the function $F(z)=\nabla f(z),\,z\in\mathbb{C}\backslash\Gamma$ (defined almost everywhere) is in $L^{1}_{loc}(\mathbb{C})$, we have, in the sense of distributions, $\nabla f=F$.
@@
Let $\varphi$ be a $C_{c}^{\infty}$-function in $\mathbb{R}$ with support included in [1,1][-1,1]. As usual, we define the approximation of identity $\varphi_{\varepsilon}(x)=\frac{1}{\varepsilon}\varphi(\frac{x}{\varepsilon})$. From this we can define the approximation of identity in two variables
@@
We may assume without loss of generality that $f$ has compact support, and then the function $f*\psi_{\varepsilon}$ is a $C^{\infty}$-function with compact support and converges uniformly in $\mathbb{C}$ towards $f$. From this we deduce that
@@
We divide the left-hand side into two parts:
Let $\Gamma_{\varepsilon}$ be the set of points at a distance less than $2\varepsilon$ from $\Gamma$. Then
@@
because $\int_{[-\varepsilon,\varepsilon]}\varphi_{\varepsilon}^{\prime}(t_{1})dt_{1}=0$. Now $|f(x-t_{1},y-t_{2})-f(x,y-t_{2})|\leq C\varepsilon^{\alpha}$ and $|\varphi_{\varepsilon}^{\prime}(t_{1})|\leq C\varepsilon^{-2}$. Since we integrate on an interval of length $2\varepsilon$, the absolute value of the entire integral is bounded by $C\varepsilon^{\alpha-1}$.
@@
But we know $|K\Gamma_{\varepsilon}|\varepsilon^{h(\Gamma)-2}\leq C$; we conclude that
@@
Based on Proposition 3.7, a sufficient condition for a positive answer to the question is the hypotheses of Theorem 3.5 and $s>\frac{1}{p}+(h(\Gamma)-1)$. Notice that this condition need not be compatible with the assumptions of the theorem, in particular with the fact that $\omega\in A_{p}$: we indeed need
$\frac{1}{p}+h-1<1+\frac{1-h}{p},$
which is equivalent to
$p>\frac{h}{2-h},$
and this condition is a new constraint if $h>4/3$.
3.4 Almost-Dirichlet principle
We start by recalling the classical Dirichlet principle. If $\Omega$ is a Jordan domain with $\partial\Omega=\Gamma$ and $f$ is a continuous function on $\Gamma$ having a continuous extension $F:\overline{\Omega}\to\mathbb{C}$ that is in $C^{1}(\Omega)$. then, if $u$ stands for the harmonic extension of $f$ to $\Omega$, we have